# An interpretable generative multimodal neuroimaging-genomics framework for decoding Alzheimer’s disease

### v1

@@ 2.5.1 Testing and performance evaluation
The best model, in terms of validation accuracy, obtained from the CV procedure of training Step 2 was considered for the testing phase for the two classification tasks presented in Section 2.3. The testing set retained from the AD and CN study cohort splitting was used for classification Task 1, AD detection. The full MCI cohort, never considered during the training phase, hence kept completely unseen by the model, was considered as a testing set for the classification Task 2, MCI conversion prediction. The framework was finally evaluated in terms of accuracy (ACC), precision (PRE), recall (REC), F1111 score, Area Under Precision Recall Curve (AUPRC), and Area Under the Curve (AUC) on both testing sets for the two classification tasks.
Five independent runs reshuffling the train and test sets were performed for probing the generalization capabilities of the model, and the average test performance over the different runs was retained. All the training and testing described were performed in Python, specifically relying on the Pytorch library [52].
@@ 3 Post-hoc interpretability analysis
The post-hoc interpretability analysis was performed relying on Integrated Gradients (IG) [53], aiming at obtaining attribution maps describing the relevance score associated with each input feature on the different input channels. IG maps were derived for the subjects in the testing set of the best model over the five generalization runs for both the classification tasks (Task 1: AD detection and Task 2: MCI prediction conversion, which coincides with the full MCI cohort). A brief introduction of the IG method and its properties followed by the attribution analysis performed for each input channel will be given in the following sections.
3.1 Integrated Gradients (IG)
Integrated Gradients (IG) [53] is an interpretability method allowing to find the relevance of the input features with respect to the prediction of the model without requiring any modification to the network under analysis. It is a Baseline Attribution Method (BAM) [54]. To obtain feature attributions, IG computes the path integral of the gradients along the straightline path from a given baseline $x^{\prime}$ to the input $x$ [53] following the equation
@@ 3.1 Integrated Gradients (IG)
where $F$ is the model and $\alpha$ is a factor ranging in [$0,1$], determining the steps along the straightline path from $x$ to $x^{\prime}$.
This method is gaining high popularity among the post-hoc interpretability methods since it satisfies three important axioms: (i) Sensitivity, for which a null score is given to the input features which do not contribute to the prediction; (ii) Implementation invariance, hence given a certain input, the derived attributions for two functionally equivalent networks are as well equivalent; and (iii) Completeness, which ensures that the attributions approximately add up to the difference between the input and the baseline prediction scores [53].
Following the definition, IG attribution scores can assume both positive and negative values. The higher or lower the value of a given feature is, the higher its contribution to the prediction outcome is, either with a positive or negative effect.
It is evident how the baseline choice impacts on the obtained attribution maps. Some recent studies have investigated this issue in order to raise awareness among the users and show the impact of different baselines on the IG scores obtained for the same input and model [55, 56, 57]. Of note, the main requirement for the baseline is to represent a neutral input for the model under investigation. In other words, it should produce a zero score or, equivalently, around $0.5$ prediction probability for both classes resulting from the activation function of the final classification layer. For the proposed framework, a triplet of neutral baselines, one for each input modality, was selected after exhaustive empirical research, defined as follows: (i) $3$D matrix of zeros for sMRI; (ii) 1D vector of zeros for rs-fMRI; and (iii) 1D vector of Gaussian noise for SNPs, which allowed on average to obtain random classification scores.
3.2 IG feature attribution analysis
IG allows obtaining attribution maps for each input and channel, which lay in the same input space. In our study IG was applied on the testing set of each classification task, resulting in a triplet of attribution maps representing sMRI, rs-fMRI, and SNPs feature attributions lying in the same space as the input data, for each subject and task. The different maps will be referred to as sMRI-IG, fMRI-IG, and SNPs-IG in what follows.
@@ 3.2.1 Neuroimaging
In order to retain the most relevant features, an initial sMRI-IG map thresholding was carried on, selecting only the attribution values exceeding in absolute value the $99.5^{th}$ percentile of the respective IG values distribution voxelwise. The average sMRI-IG values for $55$ cortical and subcortical brain regions based on the Harvard-Oxford atlas [58] were then extracted for all the subjects in the testing sets of both classification tasks. Only the regions where at least $99\%$ of the subjects had average values exceeding the percentile threshold were kept for subsequent analysis.
Moving to the fMRI-IG, only the connections whose attribution scores were in absolute value above the $98^{th}$ percentile of the respective IG value distribution connection-wise were considered for further analysis. Different thresholds were chosen between the IG maps due to the different input sizes.
A statistical group-based analysis was then performed considering the attribution maps derived for each class, namely CN, MCInc, MCIc, and AD. Shapiro-Wilk test of normality [59] was initially performed on the sMRI-IG and fMRI-IG derived features, revealing non-normally distributed features. The non-parametric Kruskal-Wallis test [60] was hence performed to check the statistical difference across the four classes considered in each feature derived either from rs-fMRI (most relevant connections resulting from fMRI-IG) or sMRI (most relevant brain regions derived from the sMRI-IG). When significance was found, the pairwise Wilcoxon Rank Sum Test [61] was performed between all the possible couples of the four considered groups (AD-CN, AD-MCIc, AD-MCInc, CN-MCIc, CN-MCInc, MCIc-MCInc) and adjusted for multiple comparisons with Bonferroni correction. All the statistical analyses were performed in R.
@@ 3.2.2 Genetics
Only the SNPs with attribution greater than the $65^{th}$ percentile of the SNPs-IG distribution, hence those having positive attribution, were kept. The Ensembl Variant Effect Predictor (VEP) tool [62] was used to annotate the selected SNPs in their corresponding genes. Web interface of VEP (https://useast.ensembl.org/Tools/VEP) and default settings were used to run the analysis. Selected settings included finding the co-located known variants and 1000 Genomes global minor allele frequency as frequency data for co-located variants parameters. No filtering was applied to the analysis to not exclude relevant biological data. Enrichment analysis was then performed using the clusterProfiler package in R [63] to derive the most important biological processes in which the annotated genes were represented. A background of $16714$ genes was selected by annotating the whole SNPs-set of the GWAS study used for feature selection of the input data. The enrichment results were corrected with the Benjamini-Hochberg procedure. The SNPs annotation, as well as the enrichment analysis, was computed only on the classes of patients, namely the AD, MCInc, and MCIc individuals.
@@ 4 Results
Classification performance and IG attribution analysis results will be firstly presented separately for classification Task 1 and Task 2. An overview of the overall group-based analysis will then follow.
@@ 4.1 Task 1: AD detection
The proposed framework reached $0.926\pm 0.02$ in ACC, $0.910\pm 0.05$ in PRE, $0.876\pm 0.03$ in REC, $0.891\pm 0.03$ in F1111 score, $0.829\pm 0.03$ in AUPRC, and $0.960\pm 0.01$ in AUC for the differentiation between AD and CN subjects. Of note, the best model over the five generalization runs reached $0.964$ in ACC, $0.984$ in PRE, $0.909$ in REC, $0.945$ in F1111, $0.876$ in AUPRC, and $0.970$ in AUC.
@@ 4.1.1 Feature relevance
Figure 4: Overview of IG attribution maps for the classification Task 1, AD detection. All the IG maps are presented for the correctly classified CN and AD subjects in the testing set. A. Average sMRI-IG maps, thresholded to retain IG scores exceeding the $99.5^{th}$ percentile and overlaid to MNI152152152152 template; B. Average fMRI-IG derived connectograms, thresholded to retain the connections with an associated IG score over the $98^{th}$ percentile; C. Average SNPs-IG scores highlighting the SNPs with an associated positive IG score exceeding the $65^{th}$ percentile, SNPs are grouped by chromosome.
@@
Fig. 4 shows the average IG maps obtained for the AD detection task and for each input channel, averaged over the correctly classified subjects per class (test set). The jet colorbar was used to highlight the attribution scores, where red and blue correspond to positive and negative attribution values, respectively. In detail, Fig. 4A shows the average sMRI-IG maps for the correctly classified CN and AD patients, overlaid onto the MNI152152152152 template (1.5 mm). A thresholding was applied in order to visualize only the relevance scores above the $99.5^{th}$ percentile of the relevance distribution, followed by a Gaussian smoothing with FWHM = 3 mm for visualization purposes. It is evident that the subcortical regions, in particular the hippocampus, are associated with high IG scores in absolute value, with the highest positive (negative) values for CN (AD). Cortical regions generally showed almost null relevance for both the CN and the AD-derived sMRI-IG maps. Of note, higher absolute values were found for the AD maps compared with the CN ones.
Moving to the fMRI-IG qualitative analysis, Fig. 4B shows the average connectograms for CN and AD subjects, including only the connections above the $98^{th}$ percentile of the relevance distribution. The ICs-brain region correspondences are detailed in Supplementary Table $2$.
The most relevant RSNs were the primary information processing-related networks (SM and VI) followed by multisensory integration networks (CC and DM), and cerebellum (CB), showing high relevance for both the CN and the AD-derived fMRI-IG. In particular, the $53$% of the most relevant connections for the CN group belong to the SM network, as opposed to the $38$% found for the AD-derived fMRI-IG maps. On the other hand, the VI was highly involved for AD subjects, with the $28$% of relevant connections belonging to this RSN, differently from the CN where only $2$ VI ICs resulted as relevant. Of note, a total of $15$ ICs were marked as relevant for both the CN and the AD-derived fMRI-IG, but with an opposite sign.
More in detail of the relevant connections between the different ICs, the CN subjects showed positive relevant intra-network connections in the CB and SM RSNs, with particularly high scores for the connections involving the post/paracentral and parietal gyri (ICs $9,10,11,12,13$). Inter-network positively relevant connections were also found between the SM and the CB, and the SM and the CC RSNs. Only two negative IG scores were instead retrieved, related to an intra-network connection in the VI network and to an inter-network connection between CC and DM.
@@
A similar pattern was found for the AD fMRI-IG relevant connections but with generally opposite IG-associated scores. In detail, negative relevance was mainly found for both the intra- and inter-network connections encompassing the SM and CB RSNs, with an intra-connection in SM (ICs $11-12$) showing the highest negative relevance between the same ICs highlighted with the opposite sign in the CN. On the contrary, positive relevance was recorded for the connections between multiple VI ICs, in particular involving right middle occipital gyrus (IC $21$) with cuneus (IC $20$), inferior occipital gyrus (IC $23$), and middle temporal gyrus (IC $25$) as well as the inter-network connections VI-DM and CC-DM, with the latter being the most relevant one.
@@
Finally, concerning the SNPs-IG qualitative analysis, Fig. 4C presents a Manhattan plot including only the SNPs with an attributed positive IG score higher than the $65^{th}$ percentile, grouped by chromosome. The y-axis reports the associated IG score. A complementary trend between the IG values associated with the SNPs was found between the two classes. This was particularly evident for Chr $1$, $7$, $11$, $14$, $15$ and $19$. A generally high involvement of the Chr $7$, $8$, $11$, $14$, and $19$ was present for both the AD and CN, with the majority of SNPs selected in these Chr showing high IG scores, either positively or negatively. Of interest, the most relevant SNPs, with positive IG were found in Chr $4$, $7$, and $8$ for the CN and in Chr $12$ and $14$ for the AD. On the contrary, the most negatively relevant SNPs for CN were found in Chr $14$ and $19$, as opposed to AD where they were mainly in Chr $1$, $6$, and $7$.
@@ 4.2 Task 2: MCI conversion prediction
The MCI prediction task was performed by testing the full MCI cohort on the best model obtained for the AD detection task, for each one of the five independent runs. This allowed to assess its viability in predicting MCI conversion to AD, by correctly stratifying MCIc and MCInc subjects, without being trained for the specific task. This procedure allowed to reach an ACC of $0.711\pm 0.01$, a PRE of $0.558\pm 0.03$, a REC of $0.610\pm 0.03$, $0.581\pm 0.01$ in F1111 score, $0.470\pm 0.02$ in AUPRC,and finally an AUC of $0.755\pm 0.01$ for the differentiation between MCIc and MCInc. Regarding the performance retrieved using the best model trained on the classification Task 1, we achieved $0.711$ in ACC, $0.612$ in PRE, $0.550$ in REC, $0.579$ in F1111 score, $0.505$ in AUPRC, and $0.766$ in AUC.
@@ 4.2.1 Feature relevance
Figure 5: Overview of IG attribution maps for the classification Task 2, MCI conversion prediction. All the IG maps are presented for the correctly classified MCIc and MCInc subjects in the testing set. A. Average sMRI-IG maps, thresholded to retain IG scores exceeding the $99.5^{th}$ percentile overlaid to MNI152152152152 template; B. Average fMRI-IG derived connectograms, thresholded to retain the connection with an associated IG score over the $98^{th}$ percentile; C. Average SNPs-IG scores highlighting the SNPs with an associated positive IG score exceeding the $65^{th}$ percentile, grouped by chromosome.
@@
In parallel with the results shown for Task 1, Fig. 5 shows a summary of the IG results obtained for the MCI conversion prediction task. The same visualization techniques presented for Task 1 were applied also for this figure. The maps were averaged over the correctly classified subjects for each class, the jet colorbar was used to highlight the attribution scores, and thresholding was applied to retain only the values above the $99.5^{th}$, $98^{th}$ and $65^{th}$ percentile of the relevance distribution, respectively for the sMRI-IG, fMRI-IG, and the SNPs-IG. The sMRI-IG maps were smoothed and overlaid onto the MNI152152152152 template. In detail, Fig. 5A shows the average sMRI-IG maps for the correctly classified MCIc and MCInc patients. Similarly to what was shown for Task 1, the subcortical regions held the highest relevance, with the hippocampus being clearly highlighted as the most important region with positive attributions for MCInc and negative values for MCIc. Cortical regions did not show particular relevance, except for some temporal and occipital areas, which however showed low IG scores.
Moving to the fMRI-IG, Fig. 5B shows the averaged connectograms. In accordance with Task 1, the same five RSNs resulted as the most relevant also for the MCI prediction conversion task, namely the SM, VI, CC, DM and CB. The MCInc fMRI-IG showed more involvement of VI ($25$% of the relevant ICs), compared with the MCIc one ($14$% of the relevant ICs). On the contrary, the MCIc IG connectome showed a higher number of ICs belonging to the CC network ($4$ ICs) compared to the two found for the MCInc one. Of note, the two connectograms had $11$ common connections with opposite trends.
@@
More in details of the relevant connections between the different ICs, of interest, the most relevant connections between the ICs in the SM were exactly the same for MCIc and MCInc, with opposite signs, and mainly involved the post/paracentral and parietal gyri (ICs $9,10,11,12,13$), as for Task 1. Moreover, the fMRI-IG for the MCInc subjects showed positive relevance also for the intra-network connections in the CB, and for the inter-network connections involving CB-SM and CC-SM. On the other hand, negative relevance was found for the intra-network connections in the VI, mainly involving occipital, lingual, calcarine, middle temporal gyri, and cuneus areas (ICs $17-24$, $18-20$, $21-25$, $23-24$), and for the inter-network connections between VI and DM, and CC and DM. A slightly different pattern was found for the MCIc fMRI-IG, whose relevant connections generally had a negative IG-associated score. In detail, intra-network connections with negative scores were found for SM and CB, as already presented, while inter-network negative connections were depicted for VI-CC, VI-DM, and SM-DM. Of interest, negative connections between CC and SM were also retrieved, never reported for the other classes in the study. Positive attribution was found only for an intra-network connection in the VI (ICs $23-24$) and a connection between CC and DM (ICs $28-45$).
@@
Finally, concerning the SNPs-IG, Fig. 5C presents a Manhattan plot highlighting the SNPs with an attributed positive IG score higher than the $65^{th}$ percentile. As expected and as already seen for the other modalities, a complementary trend between the IG values associated with the SNPs was found for the two classes, with positive weights being associated with the MCInc and negative weights associated with the MCIc for the same SNPs and vice-versa. This was particularly evident for Chr $1$, $2$, $7$, $14$, $15$ and $17$. A generally high involvement of the Chr $11$ and $17$ was present for both the MCIc and MCInc, with the majority of SNPs showing high IG scores. Of interest, the most relevant SNPs, with positive IG were found in Chr $17$ for the MCInc and in Chr $15$ for the MCIc. On the contrary, the most negatively relevant SNPs for MCInc were found in Chr $14$, $15$, and $17$, as opposed to the MCIc where they were mainly in Chr $1$, $2$, and $7$.
Figure 6: Overview of the neuroimaging statistical analysis. A. Boxplots showing the CN, AD, MCInc, and MCIc correctly classified subjects distribution in the most relevant cortical and subcortical brain regions, resulting from the sMRI-IG; B. sMRI-IG Wilcoxon test results for each group comparison and each brain region, divided into cortical and subcortical regions; C. fMRI-IG Wilcoxon test results for each group comparison for each relevant brain connection divided into extra- (upper) and intra-network (lower) connectivity. The p-values for B. and C. are reported in negative logarithmic scale, dark red the most significant. White cells represent no statistical significance.

### v3

@@ 2.4.1 Testing and performance evaluation
The best model, in terms of validation accuracy, obtained from the CV procedure of training Step 2 was considered for the testing phase for the two classification tasks. The full MCI cohort, never considered during the training phase, hence kept completely unseen by the model, was considered as a testing set for the classification Task 2, MCI conversion prediction. The framework was finally evaluated in terms of accuracy (ACC), precision (PRE), and recall (REC) on both testing sets for the two classification tasks. Five independent runs reshuffling the train and test sets were performed for probing the generalization capabilities of the model, and the average test performance over the different runs was retained.
2.5 Comparison with baseline approaches
Two baseline methods with two different imputation strategies were implemented for the sake of comparison with the proposed framework in each Task. In detail, an ensemble SVM and an ensemble random forest classifiers with mean and zero imputation strategies were developed on the dataset considered in this study. Similar to the proposed framework, the cohort composed of CN and AD was used to train and test both ensemble models, and then Task 2 (MCI cohort) was employed only in the testing phase as an external test. Five independent runs, as in the proposed method, were performed to assess the ability to generalize on different testing splits.
@@ 3 Post-hoc interpretability analysis
The post-hoc interpretability analysis was performed relying on Integrated Gradients (IG) [61], aiming at obtaining attribution maps describing the relevance of each input feature on the different input channels with respect to the outcome of the model. IG is a baseline attribution method, in the sense that, to obtain feature attributions, IG computes the path integral of the gradients along the straightline path from a given baseline $x^{\prime}$ to the input $x$ [61] following the equation
@@
where $F$ is the model and $\alpha$ is a factor ranging in [$0,1$], determining the step size along the straightline path from $x$ to $x^{\prime}$. For the proposed model, a triplet of neutral baselines, one for each input data, was selected as follows: $3$D zero matrix, 1D zero vector, and 1D Gaussian noise vector for sMRI, rs-fMRI, and SNPs, respectively.
In our study IG was applied on the testing set of each classification task, resulting in a triplet of attribution maps for sMRI, rs-fMRI, and SNPs lying in the same space as the input data, for each subject and task. The different maps will be referred to as sMRI-IG, fMRI-IG, and SNPs-IG in what follows.
@@ 3.0.1 Neuroimaging
In order to retain the most relevant features, an initial sMRI-IG map thresholding was applied, selecting only the attribution values exceeding in absolute value the $99.5^{th}$ percentile of the respective IG values distribution voxelwise. The average sMRI-IG values for $55$ cortical and subcortical brain regions based on the Harvard-Oxford atlas [62] were then extracted for all the subjects in the testing sets of both classification tasks. Only the regions where at least $99\%$ of the subjects had average values exceeding the percentile threshold were kept for subsequent analysis. For fMRI-IG, only the connections whose attribution scores were in absolute value above the $98^{th}$ percentile of the respective IG value distribution connection-wise were considered for further analysis. Different thresholds were chosen for the IG maps due to the different input sizes.
A statistical group-based analysis was then performed considering the attribution maps derived for each class, namely CN, MCInc, MCIc, and AD. Shapiro-Wilk test of normality was initially performed on the sMRI-IG and fMRI-IG derived features, revealing non-normally distributed values. The non-parametric Kruskal-Wallis test was hence performed to check the statistical difference across the four classes considered in each feature, derived either from rs-fMRI (most relevant connections resulting from fMRI-IG) or sMRI (most relevant brain regions derived from the sMRI-IG). When significance was found, the pairwise Wilcoxon Rank Sum Test was performed between all the possible pairs of the four considered groups (AD-CN, AD-MCIc, AD-MCInc, CN-MCIc, CN-MCInc, MCIc-MCInc) and adjusted for multiple comparisons with Bonferroni correction.
@@ 3.0.2 Genetics
Only the SNPs with attribution greater than the $65^{th}$ percentile in absolute terms of the SNPs-IG distribution, and those having positive attribution, were kept. The Ensemble Variant Effect Predictor (VEP) tool [63] was used to annotate the selected SNPs in their corresponding genes, considering the default settings. Enrichment analysis was then performed using the Gene Ontology enrichment analysis tool [64, 65, 66] to derive the most important biological processes in which the annotated genes were involved. A background of $334$ genes was generated by annotating the $565$ SNPs-set used as input data for the model. Bonferroni correction for multiple testing was performed, and both corrected and no corrected p-values were reported. The SNPs annotation, as well as the enrichment analysis, was computed only on the classes of patients (AD, MCInc, and MCIc).
@@ 4.1 Task 1: AD detection
The proposed framework reached $0.926\pm 0.02$ in ACC, $0.910\pm 0.05$ in PRE, and $0.876\pm 0.03$ in REC for the differentiation between AD and CN subjects. Of note, the best model over the five generalization runs reached $0.964$ in ACC, $0.984$ in PRE, and $0.909$ in REC.
Regarding the comparison with the baseline approaches for Task 1, the SVM with mean and zero imputation strategy achieved average ACCs of $0.699\pm 0.01$ and $0.689\pm 0.02$, respectively, while the random forest methods with still mean and zero missing data management approaches obtained average ACCs of $0.663\pm 0.01$ and $0.663\pm 0.01$, respectively, hence highlighting how our proposed framework outperformed simplest approaches. More information about the comparison with the baseline methods is reported in the supplementary materials "Results" section, "Classification comparison with baseline approaches" subsection.
@@ 4.1.1 Feature relevance
Figure 4: IG attribution maps for the classification Task 1. All the IG maps are presented for the correctly classified CN and AD subjects in the testing set. A. Average sMRI-IG maps, thresholded to retain IG scores exceeding the $99.5^{th}$ percentile; B. Average fMRI-IG derived connectograms, keeping only the connections over the $98^{th}$ percentile; C. Average SNPs-IG scores highlighting the SNPs with an associated IG score exceeding the $65^{th}$ percentile in absolute terms and grouped by chromosome.
@@
Figure 4 shows the average IG maps obtained for the AD detection task and for each input channel, averaged over the correctly classified subjects per class (test set). The jet colorbar was used to highlight the attribution scores, where red and blue correspond to positive and negative attribution values, respectively. In detail, Figure 4A shows the average sMRI-IG maps for the correctly classified CN and AD patients, overlaid onto the MNI152152152152 template (1.5 mm). It is evident that the subcortical regions, in particular the hippocampus, are associated with high IG scores in absolute value, with the highest positive (negative) values for CN (AD). Cortical regions generally showed almost null relevance for both the CN and the AD-derived sMRI-IG maps. Of note, higher absolute values were found for the AD maps compared with the CN ones.
@@
Moving to the fMRI-IG qualitative analysis, Figure 4B shows the average connectograms for CN and AD subjects, including only the connections above the $98^{th}$ percentile of the relevance distribution. The ICs-brain region correspondences are detailed in supplementary Table 3. The most relevant RSNs were the primary information processing-related networks (SM and VI) followed by multisensory integration networks (CC and DM), and cerebellum (CB), showing high relevance for both the CN and the AD-derived fMRI-IG. In particular, the $53$% of the most relevant connections for the CN group belong to the SM network, while a percentage of $38$% was found for the AD-derived fMRI-IG maps. On the other hand, the VI was highly involved for AD subjects, with the $28$% of relevant connections belonging to this RSN, differently from the CN where only $2$ VI ICs resulted as relevant. A total of $15$ ICs were marked as relevant for both the CN and the AD-derived fMRI-IG, but with an opposite sign. More in detail of the relevant connections between the different ICs, the CN subjects showed positive relevant intra-network connections in the CB and SM RSNs, with particularly high scores for the connections involving the post/paracentral and parietal gyri (ICs $9,10,11,12,13$). Inter-network positively relevant connections were also found between the SM and the CB, and the SM and the CC RSNs. Only two negative IG scores were instead retrieved, related to an intra-network connection in the VI network and to an inter-network connection between CC and DM. A similar pattern was found for the AD fMRI-IG relevant connections but with generally opposite IG-associated scores. In detail, negative relevance was mainly found for both the intra- and inter-network connections encompassing the SM and CB RSNs, with an intra-connection in SM (ICs 11-12) showing the highest negative relevance between the same ICs highlighted with the opposite sign in the CN. On the contrary, positive relevance was recorded for the connections between multiple VI ICs, in particular involving right middle occipital gyrus (IC $21$) with cuneus (IC $20$), inferior occipital gyrus (IC $23$), and middle temporal gyrus (IC $25$) as well as the inter-network connections VI-DM and CC-DM, with the latter being the most relevant one.
@@
Finally, concerning the SNPs-IG qualitative analysis, Figure 4C presents a Manhattan plot including the SNPs with an IG score higher than the $65^{th}$ percentile in absolute terms, grouped by chromosome. The y-axis reports the associated IG score. A complementary trend between the IG values associated with the SNPs was found between the two classes. This was particularly evident for Chr $1$, $7$, $11$, $14$, $15$, and $19$. A generally high involvement of the Chr $7$, $8$, $11$, $14$, and $19$ was present for both the AD and CN, with the majority of SNPs selected in these Chr showing high IG scores, either positively or negatively. Of interest, the most relevant SNPs, with positive IG were found in Chr $4$, $7$, and $8$ for the CN and in Chr $12$ and $14$ for the AD. On the contrary, the most negatively relevant SNPs for CN were found in Chr $14$ and $19$, as opposed to AD where they were mainly in Chr $1$, $6$, and $7$.
@@ 4.2 Task 2: MCI conversion prediction
The MCI prediction task was performed by testing the full MCI cohort on the best model obtained for the AD detection task, for each one of the five independent runs. This allowed to assess its viability in predicting MCI conversion to AD, by correctly stratifying MCIc and MCInc subjects, without being trained for the specific task. This procedure allowed to reach an ACC of $0.711\pm 0.01$, a PRE of $0.558\pm 0.03$, and a REC of $0.610\pm 0.03$ for the differentiation between MCIc and MCInc. Regarding the performance obtained using the best model trained on Task 1, we achieved $0.711$ in ACC, $0.612$ in PRE, and $0.550$ in REC.
The baseline approaches were tested under the same training strategy as the proposed model also for Task 2, where the SVM with mean and zero imputation strategy obtained ACCs of $0.690\pm 0.01$ and $0.688\pm 0.01$, respectively, and ACCs of $0.698\pm 0.01$ and $0.694\pm 0.01$ for random forest with mean and zero strategies, respectively. Additional details related to the comparison are reported in supplementary materials "Results" section, "Classification comparison with baseline approaches" subsection.
@@ 4.2.1 Feature relevance
Figure 5: IG attribution maps for the classification Task 2. All the IG maps are presented for the correctly classified MCIc and MCInc subjects in the testing set. A. Average sMRI-IG maps, thresholded to retain IG scores exceeding the $99.5^{th}$ percentile; B. Average fMRI-IG derived connectograms, keeping only the connection over the $98^{th}$ percentile; C. Average SNPs-IG scores highlighting the SNPs with an associated IG score exceeding the $65^{th}$ percentile in absolute terms and grouped by chromosome.
@@
In parallel with the results shown for Task 1, Figure 5 shows a summary of the IG results obtained for the MCI conversion prediction task. The same visualization techniques presented for Task 1 were applied also for this figure. The maps were averaged over the correctly classified subjects for each class, and thresholding was applied to retain only the values above the $99.5^{th}$, $98^{th}$, and $65^{th}$ percentile of the relevance distribution, respectively for the sMRI-IG, fMRI-IG, and the SNPs-IG. In detail, Figure 5A shows the average sMRI-IG maps for the correctly classified MCIc and MCInc patients. Similarly to what was shown for Task 1, the subcortical regions held the highest relevance, with the hippocampus being clearly highlighted as the most important region with positive attributions for MCInc and negative values for MCIc. Cortical regions did not show particular relevance, except for some temporal and occipital areas, which however showed low IG scores.
@@
Moving to the fMRI-IG, Figure 5B shows the averaged connectograms. In accordance with Task 1, the same five RSNs resulted as the most relevant also for the MCI prediction conversion task, namely the SM, VI, CC, DM and CB. The MCInc fMRI-IG showed more involvement of VI ($25$% of the relevant ICs), compared with the MCIc one ($14$% of the relevant ICs). On the contrary, the MCIc IG connectome showed a higher number of ICs belonging to the CC network ($4$ ICs) compared to the two found for the MCInc one. The two connectograms had $11$ common connections with opposite trends. More in details of the relevant connections between the different ICs, of interest, the most relevant connections between the ICs in the SM were exactly the same for MCIc and MCInc, with opposite signs, and mainly involved the post/paracentral and parietal gyri (ICs $9,10,11,12,13$), as for Task 1. Moreover, the fMRI-IG for the MCInc subjects showed positive relevance also for the intra-network connections in the CB, and for the inter-network connections involving CB-SM and CC-SM. On the other hand, negative relevance was found for the intra-network connections in the VI, mainly involving occipital, lingual, calcarine, middle temporal gyri, and cuneus areas (ICs 17-24, 18-20, 21-25, 23-24), and for the inter-network connections between VI and DM, and CC and DM. A slightly different pattern was found for the MCIc fMRI-IG, whose relevant connections generally had a negative IG-associated score. In detail, intra-network connections with negative scores were found for SM and CB, as already presented, while inter-network negative connections were depicted for VI-CC, VI-DM, and SM-DM. Of interest, negative connections between CC and SM were also retrieved, never reported for the other classes in the study. Positive attribution was found only for an intra-network connection in the VI (ICs 23-24) and a connection between CC and DM (ICs 28-45).
@@
Finally, concerning the SNPs-IG, Figure 5C presents a Manhattan plot highlighting the SNPs with an IG score higher than the $65^{th}$ percentile in absolute terms. As expected and as already seen for the other modalities, a complementary trend between the IG values associated with the SNPs was found for the two classes, with positive weights being associated with the MCInc and negative weights associated with the MCIc for the same SNPs and vice-versa. This was particularly evident for Chr $1$, $2$, $7$, $14$, $15$ and $17$. A generally high involvement of the Chr $11$ and $17$ was present for both the MCIc and MCInc, with the majority of SNPs showing high IG scores. The most relevant SNPs, with positive IG were found in Chr $17$ for the MCInc and in Chr $15$ for the MCIc. On the contrary, the most negatively relevant SNPs for MCInc were found in Chr $14$, $15$, and $17$, as opposed to the MCIc where they were mainly in Chr $1$, $2$, and $7$.